OpenStudy (anonymous):

What is the value of expression? 6^−7 • 6^4

OpenStudy (anonymous):

\[\frac{ 1 }{ 216}\] (Decimal: 0.00463)

OpenStudy (anonymous):

I might be wrong idk

rishavraj (rishavraj):

\[a^x \times a^y = a^{x+y}\]